# Sequential model confidence sets

### v1

@@ 2.1 Discussion of the different sets of superior models
Studying different notions of superior models is important, since, depending on the particular situation, one might interpret the term “best” differently. In this section, we discuss the hypotheses which correspond to the sequences of superior models given at (1), (2), and (3), respectively, in more detail, and give concrete examples for each of them.
@@
If it is reasonable to assume that some models outperform all other models at all time points (in terms of the conditional expected score differences), then we should try to find estimators for $(\mathcal{M}^{\textrm{s},\star}_{t})_{t\in\mathbb{N}}$. We refer to this assumption as the strong hypothesis. The strong hypothesis particularly applies if we assume independent and identically distributed scores. Then, one typically studies the unconditional expected score differences $\mathbb{E}(d_{ij,t})$ independent of $t\in\mathbb{N}$, as it is done by Diebold and Mariano (1995) or Hansen et al. (2011), who assume stationary performances of the models. Another important instance where the strong hypothesis applies is if the models have nested information sets, as it is e.g. the case in Example 2.2 if the different sets of covariates are nested. Then, the regression models with many covariates typically estimate the given data more accurately, and the AIC or BIC take the corresponding complexities into account by a penalty on the regression dimension. In Section 5.1, we construct SMCSs for different Covid-19 related deaths where the assumption of nested information sets seems plausible.
@@
On the other hand, if we assume that some models have a lower conditional expected score on average rather than at all single time points, then we should target $(\mathcal{M}^{\textrm{uw},\star}_{t})_{t\in\mathbb{N}}$ under the uniformly weak hypothesis. Choe and Ramdas (2023, Section 4.4.) argue convincingly that there are many situations where testing for the strong hypothesis may be misleading. Inspired by their arguments, Example 2.4 below considers a forecaster which is ideal on most of the days and only slightly worse than the other forecasters on Sundays. Any powerful method to monitor the strongly superior models over time will exclude this forecaster from the SMCS since she is indeed not strongly superior. Nevertheless, the forecaster is superior under the uniformly weak hypothesis as she is generally better than all other forecasters and only slightly worse than some of them on Sundays.
@@
Finally, we say that the weak hypothesis applies if we expect that also the average performances might change over time. The weak hypothesis is of particular interest if we assume that the models might evolve over time and improve their outputs individually by using past observations. In Section 5.2, we study post-processing methods for wind gust predictions. The results show that in this application it is indeed beneficial to monitor the weakly superior objects as some models which are excluded from the SMCS at earlier time points are included again at later stages, either due to systematic changes in the underlying meteorological model, or due to the different adaptive behaviour of the methods. If we had conducted the study under any of the two stronger hypothesis, we could not have observed that some methods become again competitive towards the end of the test period, since, under these hypotheses, a model can not be included again in the SMCS once it is excluded.
Since the weak hypothesis is the most general, one might wonder, why we do not only consider this particular hypothesis as it is implied by the other two hypotheses anyhow? The answer to this question is given in terms of power: Whereas we impose more assumptions on the data-generating process for the former hypotheses, our methods for the latter are generally less powerful if we apply them to construct SMCSs for the strongly (uniformly weakly) superior models. However, if we construct SMCSs for the strongly (uniformly weakly) superior models even though the corresponding hypotheses do not apply, then we may expect misleading results as argued before.
@@
To conclude, we introduce the simulation settings from Section 4 to give concrete examples for each of the discussed hypotheses.
Example 2.3 (Simulation 1).
For $n=1000$, we sample $(y_{t})_{t=1}^{n}$, where $y_{t}\sim\mathcal{N}(y_{t-1},1),$ for $t=1,\dots,n$ and $y_{0}=0$. We consider $m=49$ different forecasters $\{1,\dots,m\}$ which all sequentially issue predictive distributions for $(y_{t})_{t=1}^{n}$ given by
$f_{i,t}=\mathcal{N}\left(y_{t-1}+\varepsilon_{i},1+\delta_{i}\right),\quad t=1,\dots,n,$
@@
for $(\varepsilon_{i},\delta_{i})\in\{-0.6,-0.4,-0.2,0,0.2,0.4,0.6\}^{2}$. For these forecasters with different biases and dispersion errors, there is exactly one ideal forecaster ${i_{0}}$ with $\varepsilon_{i_{0}}=\delta_{i_{0}}=0$. We assess the different forecasters with respect to the continuous ranked probability score (CRPS; Matheson and Winkler (1976)), see Appendix A. Since the CRPS is a proper scoring rule, it follows that $\mathcal{M}^{\textrm{s},\star}_{t}=\{i_{0}\}$ for all $t\in\mathbb{N}$.
Example 2.4 (Simulation 2).
We consider the same data-generating mechanism and the same forecasters as in Example 2.3 with the only exception that the forecaster ${i_{0}}$ now issues the forecasts $f_{i_{0},t}=\mathcal{N}\left(y_{t-1}+\varepsilon_{i_{0},t},1+\delta_{i_{0},t}\right)$, $t=1,\dots,n$, for
$\varepsilon_{i_{0},t}=\delta_{i_{0},t}=\begin{cases}0.3&,t\in 7\mathbb{N}\\ 0&,\textrm{else}.\end{cases}$
@@
That is, forecaster $f_{i_{0}}$ is still ideal on, say, weekdays and Saturdays, however, on Sundays, some forecasters have a smaller bias and dispersion error, and hence $\mathcal{M}^{\textrm{s},\star}_{t}=\emptyset$ for all $t\geq 7$. However, ${i_{0}}$ is uniformly weakly superior, that is $\mathcal{M}^{\textrm{uw},\star}_{t}=\{{i_{0}}\}$ for all $t\in\mathbb{N}$.
Example 2.5 (Simulation 3).
We sample i.i.d. standard normally distributed observations $(y_{t})_{t=1}^{n}$, for $n=800$, and compare $m=3$ different forecasters $\{1,2,3\}$ which issue median predictions $m_{i,t}=y_{t}+\varepsilon_{i,t}$, $i=1,2,3$, for
$\varepsilon_{1,t}=\beta,\quad\varepsilon_{2,t}=\gamma^{t},\quad\varepsilon_{3,t}=\delta t,\quad t=1,\dots,n,$ (6)
@@
for $\beta,\delta>0$ and $10<\gamma<$ 1. We call forecaster $1$ constantly biased, forecaster $2$ improving and forecaster $3$ disimproving. We assess them by the scoring function $L(m,y)=0.5(\Phi(m)-\Phi(y))$, for $m,y\in\mathbb{R}$ and $\Phi$ the cdf of the standard normal distribution, which is a consistent scoring function for the median, see Appendix A. If we choose the parameters as given in Section 4, we have $\mathcal{M}^{\textrm{w},\star}_{t}=\{3\}$ for $t\leq 153$, $\mathcal{M}^{\textrm{w},\star}_{t}=\{1\}$ for $550153<t<$ 550, and $\mathcal{M}^{\textrm{w},\star}_{t}=\{2\}$ for $t\geq 550$. That is, whereas the disproving forecaster is initially superior, the constantly biased forecaster catches up after some time until the improving forecaster becomes the best in the end.
@@ 3 Construction of sequential model confidence sets
In this section, we provide methods to construct SMCSs for the superior models given at (1), (2) and (3). Our constructions build on e-processes, confidence sequences and sequential multiple testing methods as introduced in the following.
3.1 Sequential testing methods and multiple testing
Let $(\Omega,\mathcal{F})$ be a measurable space and $\mathfrak{B}(\Omega)$ be the family of all probability measures on $(\Omega,\mathcal{F})$. A statistical (null) hypothesis $\mathcal{H}\subseteq\mathfrak{B}(\Omega)$ is a set of probability measures that are potential candidates for the true probability measure $\mathbb{P}$ governing the data generating process. We let $(\mathcal{F}_{t})_{t\in\mathbb{N}}\subseteq\mathcal{F}$ be a filtration and interpret $\mathcal{F}_{t}$ as the available information at time $t\in\mathbb{N}$. We call a process $(M_{t})_{t\in\mathbb{N}}$ a test (super-)martingale for some hypothesis $\mathcal{H}\subseteq\mathfrak{B}(\Omega)$ if it is a nonnegative (super-)martingale with $M_{0}\leq 1$ for each $\mathbb{Q}\in\mathcal{H}$. Test martingales allow for a financial interpretation in terms of betting as the accumulated wealth of a gambler after having bet a number of times against the null hypothesis (Shafer, 2021). If the null hypothesis is true, then one cannot expect to gain money over time, and thus, large values of the test martingale give evidence against the null hypothesis. More rigorously, any test supermartingale $(M_{t})_{t\in\mathbb{N}}$ for $\mathcal{H}\subseteq\mathfrak{B}(\Omega)$ satisfies
@@ 3.1 Sequential testing methods and multiple testing
The inequality at (7), referred to as Ville’s inequality (Ville, 1939), follows from the optional stopping theorem for martingales and forms a time-uniform extension of Markov’s inequality. It lays the foundation for safe anytime-valid inference by enabling the construction of time-uniform confidence sequences, sequential tests, or anytime-valid p-values amongst others (Ramdas et al., 2023).
For some problems, it is theoretically impossible to construct nontrivial test martingales (Ramdas et al., 2022). However, often one still can draw sequential inference by using so-called e-processes: An e-process for some hypothesis $\mathcal{H}\subseteq\mathfrak{B}(\Omega)$ is a nonnegative adapted stochastic process $(E_{t})_{t\in\mathbb{N}}$ with
$\mathbb{E}_{\mathbb{Q}}(E_{\tau})\leq 1$
@@
for all $\mathbb{Q}\in\mathcal{H}$ and all (possibly infinite) stopping times $\tau$. In other words, for any random stopping time $\tau$, $E_{\tau}$ is an e-variable (Vovk and Wang, 2021). Equivalently, an adapted nonnegative process $(E_{t})_{t\in\mathbb{N}}$ is an e-process for $\mathcal{H}$ if it is upper bounded by a test supermartingale for each $\mathbb{Q}\in\mathcal{H}$ and hence, any test supermartingale is also an e-process and Ville’s inequality continues to hold for e-processes (Ramdas et al., 2020).
Test martingales and e-processes allow for sequential testing. A sequential testing procedure or sequential test for some hypothesis $\mathcal{H}\subseteq\mathfrak{B}(\Omega)$ is an adapted stochastic process $(\Psi_{t})_{t\in\mathbb{N}}\subseteq\{0,1\}$, where $\Psi_{t}=1$ indicates that the hypothesis $\mathcal{H}$ is rejected at $t\in\mathbb{N}$. A sequential test $(\Psi_{t})_{t\in\mathbb{N}}$ for $\mathcal{H}\subseteq\mathfrak{B}(\Omega)$ is said to uniformly control the type-I error at level $\alpha\in(0,1)$ if
$\sup_{\mathbb{Q}\in\mathcal{H}}\mathbb{Q}(\exists t\in\mathbb{N}:\Psi_{t}=1)\leq\alpha.$
By Ville’s inequality at (7), any e-process $(E_{t})_{t\in\mathbb{N}}$ directly leads to a sequential testing procedure $(\Psi_{t})_{t\in\mathbb{N}}$ with uniform type-I error control at level $\alpha\in(0,1)$ by defining
$\Psi_{t}=\mathbbm{1}\{E_{t}\geq 1/\alpha\},\quad t\in\mathbb{N}.$ (8)
If we test more than one hypothesis simultaneously, then we have to correct for multiple testing. In the spirit of Vovk and Wang (2021), we call $m\geq 2$ e-processes $(E_{1,t})_{t\in\mathbb{N}},\dots,(E_{m,t})_{t\in\mathbb{N}}$ for the hypotheses $\mathcal{H}_{1},\dots,\mathcal{H}_{m}$ family-wise valid (FWV) if there exists a dominating family of e-processes, that is, if there exists a family $\{(E^{\mathbb{Q}}_{t})_{t\in\mathbb{N}}\mid\mathbb{Q}\in\mathfrak{B}(\Omega)\}$ such that $(E^{\mathbb{Q}}_{t})_{t\in\mathbb{N}}$ is an e-process for ${}\{\mathbb{Q}\}$ for all $\mathbb{Q}\in\mathfrak{B}(\Omega)$, and for any $i=1,\dots,m$ with $\mathbb{Q}\in\mathcal{H}_{i}$, $(E^{\mathbb{Q}}_{t})_{t\in\mathbb{N}}$ upper bounds $(E_{i,t})_{t\in\mathbb{N}}$. In Appendix C, we show that for family-wise valid e-processes, the corresponding sequential tests given at (8) time-uniformly control the family-wise error rate (FWER).
To derive SMCS for the sequence of weakly superior objects, we will make use of confidence sequences with time-uniform coverage guarantees (Howard et al., 2021; Ramdas et al., 2023). For $\alpha\in(0,1)$, a (time-uniform) $(1-\alpha)$-confidence sequence for a stochastic process $(\theta_{t})_{t\in\mathbb{N}}\subseteq\Theta$ in some space $\Theta$ is a sequence $(C_{t})_{t\in\mathbb{N}}\subseteq 2^{\Theta}$ such that
$\mathbb{Q}\left(\forall t\in\mathbb{N}:\theta_{t}\in C_{t}\right)\geq 1-\alpha,\quad\textrm{for all }\mathbb{Q}\in\mathfrak{B}(\Omega).$
3.2 SMCSs for the strongly superior models
In this section, we construct SMCSs for the sequence of strongly superior models $(\mathcal{M}^{\textrm{s},\star}_{t})_{t\in\mathbb{N}}$ given at (1) by a sequential testing procedure. Throughout this section, we assume that for any $i,j\in\mathcal{M}_{0}$, the sequence of loss differences $(d_{ij,t})_{t\in\mathbb{N}}$ is conditionally bounded with respect to some predictable sequence $(c_{ij,t})_{t\in\mathbb{N}}\subseteq\mathbb{R}_{+}$, that is $|d_{ij,t}|\leq c_{ij,t}/2$ for all $t\in\mathbb{N}$. For $i,j\in\mathcal{M}$, we consider the strong hypotheses
$\mathcal{H}^{\textrm{s}}_{ij}=\{\mathbb{Q}\in\mathfrak{B}(\Omega)\mid\mu_{ij,t}\leq 0,\forall t\in\mathbb{N}\}$
and the strong intersection hypotheses
$\mathcal{H}^{\textrm{s}}_{i\cdot}=\bigcap_{j\in\mathcal{M}_{0}}\mathcal{H}^{\textrm{s}}_{ij}=\{\mathbb{Q}\in\mathfrak{B}(\Omega)\mid\mu_{ij,t}\leq 0,\forall t\in\mathbb{N},\forall j\in\mathcal{M}_{0}\}.$
Then, for any predictable $(\lambda_{ij,t})_{t\in\mathbb{N}}$ with $0\leq\lambda_{ij,t}\leq 1/c_{ij,t}$, we have that
$M_{ij,t}=\prod_{r=1}^{t}(1+\lambda_{ij,r}(d_{ij,r}-\mu_{ij,r})),\quad t\in\mathbb{N}$
is a nonnegative martingale. Importantly, for any $\mathbb{Q}\in\mathcal{H}^{\textrm{s}}_{ij}$, $M_{ij,t}$ upper bounds
$E_{ij,t}=\prod_{r=1}^{t}(1+\lambda_{ij,r}d_{ij,r}),\quad t\in\mathbb{N},$
Hence, $(E_{ij,t})_{t\in\mathbb{N}}$ is an e-process for $\mathcal{H}^{\textrm{s}}_{ij}$, and thus also for $\mathcal{H}^{\textrm{s}}_{i\cdot}$. For $i=1,\dots,m$, we let

### v3

@@ 2.1 Discussion of the different sets of superior models
Studying different notions of superior models is important, since, depending on the particular situation, one might interpret the term “best” differently. In this subsection, we discuss the hypotheses which correspond to the sequences of superior models given at (1), (2), and (3), respectively, and give concrete examples for each of them.
@@
If it is reasonable to assume that some models outperform all other models at all time points (in terms of the conditional expected score differences), then we should try to find estimators for $\mathcal{M}^{\textrm{s},\star}$. We refer to this assumption as the strong hypothesis. The strong hypothesis particularly applies if we assume independent and identically distributed scores. Then, one typically studies the unconditional expected score differences $\mathbb{E}(d_{ij,t})$ independent of $t\in\mathbb{N}$, as it is done by Diebold and Mariano (1995) or Hansen et al. (2011), who assume stationary performances of the models. Another important instance where the strong hypothesis applies is if the models have nested information sets. In Section 5.1, we construct SMCSs for different Covid-19 related deaths where the assumption of nested information sets seems plausible.
@@
On the other hand, if we assume that some models have a lower conditional expected score on average rather than at all single time points, then we should target $\mathcal{M}^{\textrm{uw},\star}$ under the uniformly weak hypothesis. Choe and Ramdas (2023, Section 4.4.) argue convincingly that there are many situations where testing for the strong hypothesis may be misleading. Inspired by their arguments, Simulation 2 below considers a forecaster which is ideal on most days and only slightly worse than the other forecasters on Sundays. Any powerful method to monitor the strongly superior models over time would exclude this forecaster from the SMCS. Nevertheless, the forecaster is superior under the uniformly weak hypothesis.
@@
Finally, the weak hypothesis applies if we expect that the models evolve and their average relative performance might change over time. In Section 5.2, we study post-processing methods for wind gust predictions. The results show that in this application it is indeed beneficial to monitor the weakly superior objects, as some models which are excluded from the SMCS at earlier time points are included again at later stages, either due to systematic changes in the underlying meteorological model, or due to the different adaptive behaviour of the methods. If we had conducted the study under any of the two stronger hypotheses, we could not have observed that some methods become again competitive towards the end of the test period.
@@
We introduce the simulation settings of Section 4 to give concrete examples for each of the discussed hypotheses.
Simulation 1.
@@
We sample $(Y_{t})_{t=1}^{n}$, where $Y_{t}\sim\mathcal{N}(Y_{t-1},1),$ for $t=1,\dots,n$ and $Y_{0}=0$. We consider $m=49$ different forecasters $\{1,\dots,m\}$ which all sequentially issue predictive distributions for $(Y_{t})_{t=1}^{n}$ given by $f_{i,t}=\mathcal{N}(Y_{t-1}+\varepsilon_{i},1+\delta_{i})$, for $(\varepsilon_{i},\delta_{i})\in\{-0.6,-0.4,-0.2,0,0.2,0.4,0.6\}^{2}$. For these forecasters with different biases and dispersion errors, there is exactly one ideal forecaster ${i_{0}}$ with $\varepsilon_{i_{0}}=\delta_{i_{0}}=0$. We assess the different forecasters with respect to the continuous ranked probability score (CRPS; Matheson and Winkler (1976)). The CRPS is a proper scoring rule implying that the ideal forecaster has the lowest expected CRPS. According to the definition given at (1), this yields that $\mathcal{M}^{\textrm{s},\star}=\{i_{0}\}$. A brief introduction to (proper) scoring rules, including information about the CRPS, can be found in Supplement A.
Simulation 2.
@@
We consider the same data-generating mechanism and the same forecasters as in Example 1 except that, for $t\in 7\mathbb{N}$, the forecaster ${i_{0}}$ now issues the forecasts $f_{i_{0},t}=\mathcal{N}\left(Y_{t-1}+0.3,1.3\right)$. That is, forecaster $f_{i_{0}}$ is still ideal on, say, weekdays and Saturdays, however, on Sundays, some forecasters have a smaller bias and dispersion error, and hence $\mathcal{M}^{\textrm{s},\star}=\emptyset$. However, ${i_{0}}$ is uniformly weakly superior, that is $\mathcal{M}^{\textrm{uw},\star}=\{{i_{0}}\}$.
Simulation 3.
@@
We sample i.i.d. standard normally distributed observations $(Y_{t})_{t=1}^{n}$ and compare $m=3$ different forecasters $i=1,2,3$ which issue median predictions $m_{i,t}=Y_{t}+\varepsilon_{i,t}$, for $\varepsilon_{1,t}=\beta,\varepsilon_{2,t}=\gamma^{t},\varepsilon_{3,t}=\delta t,$ for $\beta,\delta>0$ and $10<\gamma<$ 1. We call forecaster $1$ constantly biased, forecaster $2$ improving and forecaster $3$ worsening. We assess them by $L(m,y)=0.5(\Phi(m)-\Phi(y))$, for $m,y\in\mathbb{R}$ and $\Phi$ the cdf of the standard normal distribution, which is a consistent scoring function for the median, see Supplement A. If we choose the parameters as given in Section 4, we have $\mathcal{M}^{\textrm{w},\star}_{t}=\{3\}$ for $t\leq 153$, $\mathcal{M}^{\textrm{w},\star}_{t}=\{1\}$ for $550153<t<$ 550, and $\mathcal{M}^{\textrm{w},\star}_{t}=\{2\}$ for $t\geq 550$. That is, whereas the worsening forecaster is initially superior, the constantly biased forecaster catches up after some time until the improving forecaster becomes the best in the end.
@@ 3 Construction of sequential model confidence sets
In this section, we provide methods to construct SMCSs for the superior models given at (1), (2) and (3). Our constructions build on e-processes, confidence sequences and sequential multiple testing methods.
3.1 Sequential testing methods
Let $(\Omega,\mathcal{F})$ be a measurable space and $\mathfrak{B}(\Omega)$ be the family of all probability measures on $(\Omega,\mathcal{F})$. Suppose that we observe random outcomes of a process at time points $t=1,2,\ldots$. A statistical (null) hypothesis $\mathcal{H}\subseteq\mathfrak{B}(\Omega)$ is a set of probability measures that are potential candidates for the true probability measure $\mathbb{P}$ governing the data generating process. For example, $\mathcal{H}$ could consist of all probability measures under which the data points are i.i.d. normally distributed with some given mean and variance.
Let $(\mathcal{F}_{t})_{t\in\mathbb{N}}$ be a filtration, that is, an increasing sequence of $\sigma$-algebras on $(\Omega,\mathcal{F})$, which are all contained in $\mathcal{F}$. We interpret $\mathcal{F}_{t}$ as the available information at time $t\in\mathbb{N}$.
A (super-)martingale with respect to a probability measure $\mathbb{Q}\in\mathfrak{B}(\Omega)$ is a sequence of $\mathbb{Q}$-integrable random variables $(X_{t})_{t\in\mathbb{N}}$, which is adapted to $(\mathcal{F}_{t})_{t\in\mathbb{N}}$ and satisfies the condition $\mathbb{E}_{\mathbb{Q}}[X_{t}|\mathcal{F}_{t-1}]=X_{t-1}$ ($\mathbb{E}_{\mathbb{Q}}[X_{t}|\mathcal{F}_{t-1}]\leq X_{t-1}$), for all $t\in\mathbb{N}$. We call a process $(M_{t})_{t\in\mathbb{N}}$ a test (super)martingale for some hypothesis $\mathcal{H}\subseteq\mathfrak{B}(\Omega)$ if it is a nonnegative (super)martingale with respect to any $\mathbb{Q}\in\mathcal{H}$, and has a starting value $M_{0}\leq 1$.
Test martingales allow to phrase statistical testing as a betting game. Indeed, the value of the test martingale can be interpreted as the accumulated wealth of a gambler after having bet a number of times against the null hypothesis (Shafer, 2021). For example, in a coin-toss game with probability of heads equal to $q\in(0,1)$, a player that bets 1\$ on heads in each round and receives 2\$ in case of success or 0\$ otherwise, cannot expect to gain money over time if the null hypothesis $\mathcal{H}_{0}:q\leq 1/2$ is true. Thus, large values of the test martingale give evidence against the null hypothesis.
Ville’s inequality (Ville, 1939) states that any test supermartingale $(M_{t})_{t\in\mathbb{N}}$ for $\mathcal{H}\subseteq\mathfrak{B}(\Omega)$ satisfies
@@ 3.1 Sequential testing methods
This lays the foundation for safe anytime-valid inference. By (5), test martingales yield valid sequential tests $(\psi_{t})_{t\in\mathbb{N}}$ by defining $\psi_{t}=\mathbbm{1}\{M_{t}\geq 1/\alpha\}$. Furthermore, Ville’s inequality enables the construction of time-uniform confidence sequences (Howard et al., 2021; Ramdas et al., 2023). For $\alpha\in(0,1)$, a $(1-\alpha)$-confidence sequence for a parameter sequence $(\theta_{t})_{t\in\mathbb{N}}$ in some space $\Theta$ is a sequence of sets $(C_{t})_{t\in\mathbb{N}}$ such that $\mathbb{Q}\left(\forall t\in\mathbb{N}:\theta_{t}\in C_{t}\right)\geq 1-\alpha$, for all $\mathbb{Q}\in\mathfrak{B}(\Omega).$ The definition of confidence sequences resembles that of confidence sets, with the important difference that confidence sequences provide time-uniform coverage guarantees. We will use confidence sequences to construct SMCSs for the sequence of weakly superior objects.
@@
It has been shown that for some hypotheses $\mathcal{H}$, it is not possible to construct non-trivial test supermartingales (Ramdas et al., 2022; Henzi and Law, 2024). However, it may still be possible to find non-trivial e-processes. An e-process for some hypothesis $\mathcal{H}\subseteq\mathfrak{B}(\Omega)$ is a nonnegative adapted stochastic process $(E_{t})_{t\in\mathbb{N}}$ with $\mathbb{E}_{\mathbb{Q}}(E_{\tau})\leq 1$ for all $\mathbb{Q}\in\mathcal{H}$ and all (possibly infinite) stopping times $\tau$. Equivalently, an adapted nonnegative process $(E_{t})_{t\in\mathbb{N}}$ is an e-process for $\mathcal{H}$ if it is upper bounded by a test supermartingale for each $\mathbb{Q}\in\mathcal{H}$. Hence, any test supermartingale is also an e-process. Importantly, Ville’s inequality continues to hold for e-processes (Ramdas et al., 2020).
When testing more than one hypothesis simultaneously with e-processes, multiple testing corrections are necessary. We discuss the relevant background in Supplement C.
3.2 SMCSs for the strongly superior and uniformly weakly superior objects
In this section, we construct SMCSs for the strongly superior models $\mathcal{M}^{\textrm{s},\star}$ given at (1) and for the uniformly weakly superior models $\mathcal{M}^{\textrm{uw},\star}$ given at (2) using e-processes and sequential testing procedures.
We define the strong hypotheses $\mathcal{H}^{\textrm{s}}_{ij}=\{\mathbb{Q}\in\mathfrak{B}(\Omega)\mid\mu_{ij,r}\leq 0,\forall t\in\mathbb{N}\}$, and the weak hypotheses $\mathcal{H}^{\textrm{uw}}_{ij}=\{\mathbb{Q}\in\mathfrak{B}(\Omega)\mid\Delta_{ij,r}\leq 0,\forall t\in\mathbb{N}\}$ for $i,j\in\mathcal{M}_{0}$. We would like to test all these pairwise hypothesis simultaneously at each time point $t\in\mathbb{N}$. For $\bullet\in\{\mathrm{s},\mathrm{uw}\}$, assume that, for any $i,j\in\mathcal{M}_{0}$, $(E_{ij,t})_{t\in\mathbb{N}}$ is an e-process for the hypothesis $\mathcal{H}^{\bullet}_{ij}$. Then, for any $i\in\mathcal{M}_{0}$, the arithmetic mean $E_{i\cdot,t}=1/(m-1)\sum_{j\neq i}E_{ij,t}$ is an e-process for the intersection hypothesis $\mathcal{H}^{\bullet}_{i\cdot}=\cap_{j\neq i}\mathcal{H}^{\bullet}_{ij}$. We adjust the e-processes $(E_{1\cdot,t})_{t\in\mathbb{N}},\dots,(E_{m\cdot,t})_{t\in\mathbb{N}}$ for multiple testing using the closure principle with the arithmetic mean as $e$-merging function, that is,